Systems and Network Management

Installing Windows in VMware on Linux

Background:

VMware is commercial software from VMware Inc. You can download the software from http://www.VMware.com/. You can apply for an evaluation licence that lasts 30 days, or buy a student licence for US $99. Our department has purchased licences for use in twenty-one computers, for version 2.0.x, with Linux as the host operating system.

Procedure:

Install the licence:

1. Obtain a VMware licence, which is a small text file

2. Create a directory .vmware in your home directory:

   $ cd
   $ mkdir .vmware

3. Put your licence file in that directory, with a file name that begins with license Install the RPM package:

1. Obtain a VMware licence, which is a small text file

2. Mount ictlab.tyict.vtc.edu.hk:/var/ftp/pub on some directory on your com-

puter (say /mnt/ftp):

   $ sudo mount nicku.org:/var/ftp/pub /mnt/ftp

3. Change to the vmware directory:

   $ cd /mnt/ftp/vmware

4. Install the VMware RPM package:

   $ sudo rpm -Uhv VMware-2.0.4-1142.i386.rpm

Build and install the kernel modules to support VMware:

1. Change to the directory where I have unpacked vmware-ws-1142-for-2.4.7.tar.gz:

   $ cd vmware-ws-1142-for-2.4.7

2. Run the runme.pl Perl script:

   $ sudo ./runme.pl

3. Accept all default choices. I may elaborate this point later; I need to install VMware again to confirm that all defaults are appropriate.

Nick Urbanik <nicku(at)nicku.org>  ver. 1.2

Installing Windows in VMware on Linux Systems and Network Management

2

Install the guest operating system (Windows):

1. Start VMware:

   $ vmware &

2. Choose “Run the Configuration Wizard”

3. Make a directory /var/vmware and make it owned by you:

   $ sudo mkdir /var/vmware
   $ sudo chown nicku.nicku /var/vmware

Please use your ldap account user name, not mine!

4. Make sure that you select this directory as the location to install the virtual machine. Do not select the default, a directory in your home directory. If you install it on a network drive, performance will be very bad, and the load on the network would be great.

5. For the size of the virtual hard disk, choose 1.5 GB. Specify that the guest operating system is Windows 2000.

6. After configuration is complete, click the power-on button. You will see the virtual machine boot, go through its power-on self-test (post).

7. Insert the Windows cdrom, and install Windows as you would on a normal non-

virtual computer Install the VMware Tools: The VMware tools are essential to allow better than 640 × 480 vga resolution.

1. Boot Windows in the virtual machine, and from the VMware menu, select Settings → VMware Tools install

2. In Windows Explorer, open the A: drive.

3. Double-click on the VMware Tools icon, and follow instructions.

4. After installation is complete, reboot the virtual computer when requested.

5. When the virtual computer reboots, let it detect your new virtual hardware.

6. Adjust the display properties to your liking. Let Nick Urbanik know if this is unclear, or I have left out any important steps.

Nick Urbanik <nicku(at)nicku.org>  ver. 1.2